# On the Burst-Covering Radius of Binary Cyclic Codes

### v1

@@ Abstract
We define and study burst-covering codes. We provide some general bounds connecting the code parameters with its burst-covering radius. We then provide stronger bounds on the burst-covering radius of cyclic codes, by employing linear-feedback shift-register (LFSR) sequences. For the case of BCH codes we prove a new bound on pattern frequencies in LFSR sequences, which is of independent interest. Using this tool, we can bound the covering-radius of binary primitive BCH codes and Melas codes. We conclude with an efficient algorithm for burst-covering cyclic codes.
@@ I Introduction
Error-correcting codes and their geometric counterparts, covering codes, have a long and rich history of research (e.g., [MacSlo78, CohHonLitLob97]). From a geometric perspective, while the former pack the space with error balls, the latter cover it. In their intersection lie perfect codes, that manage to tile the space with error balls.
Many families of codes have been studied both for the their error-correction capabilities, as well as their covering parameters. A partial list of those contains MDS codes, cyclic codes (including BCH, dual BCH, Melas and Zetterberg codes), Reed-Muller codes, and of course, perfect codes (for details, see for example [MacSlo78, CohHonLitLob97]).
However, one family of codes is conspicuously missing from this list – codes for burst errors. A $b$-burst error is an error pattern all of whose erroneous symbols are confined to a contiguous block of $b$ positions. From an error-correction perspective, such codes are motivated by the existence of bursty channels, that tend to group together erroneous positions. Burst-correcting codes have a history of research almost as long as that of error-correcting codes, starting with [Abr59, Fir59, ElsSho62]. Some of these burst-correcting codes are optimal, in the sense that they have the closest possible integer parameters to those dictated by the ball-packing bound, making them almost perfect. Among these we mention the cyclic codes of [AbdMcEOdlTil86, Abd88], and the cyclic two-dimensional codes of [SchEtz05, Construction A]. To the best of our knowledge, only one construction of perfect burst-correcting codes is known [Etz01b], for binary codes with burst length b=2b=2111There is a subtlety in the definitions of optimal and perfect burst-correcting codes: while optimal codes consider bursts cyclically, perfect codes do not.. These codes are therefore the only known burst-covering codes.
@@
Burst-covering codes do not only fill a void in the theory of binary codes, but may also find use in data storage scenarios. Many applications require the implementation of database queries whose return value is a linear combination of database items with coefficients supplied by the user. Private information retrieval (PIR) protocols [ChoGolKusSud98], partial-sum queries [ChazeRosen89] and more recently, certain machine-learning inference implementations [RamRavTam24] that are based on ideas from generalized covering radii [EliFirSch21a, EliWeiSch22], among others, all employ such queries. Traditionally, covering codes are used to answer such linear queries. This relies on the fact that any column vector can be obtained as a linear combination of at most $R$ columns of a parity-check matrix for the code, where $R$ is the code’s covering radius. This guarantees a bounded access complexity, speeding up the computation of the answer to the linear query.
@@
However, in some media types, a significant component in the time to answer a query is not only the number of items that need to be accessed, but also their spatial location (e.g., the seek time in HDDs). Such systems perform best when accessing contiguous blocks of data, as opposed to scattered random-access patterns. This resulted in a surge of interest in codes that take into account access patterns, and prioritize access in contiguous blocks [wu2021achievable, chee2024repairing, YuYuaSch25]. In the context of our work, as we will prove, the parity-check matrix $H$ of a linear $b$-burst-covering code satisfies the property that any column vector can be obtained as a linear combination of a window of at most $b$ consecutive columns of $H$. Thus, employing a burst-covering code can address locality issues in computing linear database queries.
@@
Our main contributions are as follows: We first define burst-covering codes and derive basic bounds on their parameters. We then focus on the study of the burst-covering radius of binary cyclic codes. This is motivated by the fact that cyclic codes have a rich structure, they contain some very useful code families (e.g., BCH codes), and the fact that we have almost perfect cyclic burst-correcting codes [AbdMcEOdlTil86, Abd88]. As we later show, the study of cyclic burst-covering codes is closely related to the analysis of pattern frequencies in linear-feedback shift-register (LFSR) sequences. We employ classic known bounds for the number of occurrences of subwords in these sequences, but for the relevant case of BCH codes we show that these fail to provide significant results. We thus prove a new result on pattern frequencies which is meaningful for BCH codes, and is of independent interest for the study of LFSR sequences. We use this to bound the burst-covering radius of binary BCH codes, and the closely related Melas codes. Finally, we show that our analysis naturally gives rise to an efficient covering algorithm for binary cyclic codes.
@@
The paper is organized as follows. Section II gives the necessary notation and known results to be used later. In Section III we define general linear burst-covering codes, and provide some bounds on their parameters. We then to study cyclic burst-covering codes in Section IV, and further focus on the burst-covering radius of BCH codes in Section V. An efficient covering algorithm is described in Section VI. We conclude in Section VII with a summary of the results and some open questions.
@@ II Preliminaries
Given a vector $v=(v_{0},\dots,v_{n-1})\in\mathbb{F}_{q}^{n}$, a cyclic shift of $v$ is the vector $(v_{n-1},v_{0},v_{1},\dots,v_{n-2})$. A linear code $\mathcal{C}$ is said to be cyclic if $c\in\mathcal{C}$ implies the cyclic shift of $c$ is also in $\mathcal{C}$. Denote $\mathbb{F}_{q}[X]$ the set of polynomials in the unknown $X$, with coefficients from $\mathbb{F}_{q}$. With any vector $v$ we associate the polynomial $v(X)=\sum_{i=0}^{n-1}v_{i}X^{i}$. It is well known [MacSlo78], that an $n,n-r]_{q}$ cyclic code $\mathcal{C}$ is an ideal in the ring of polynomials $\mathbb{F}_{q}[X]/(X^{n}-1)$. There exists a unique generator polynomial, $g(X)\in\mathbb{F}_{q}[X]$, $\deg(g(X))=r$, such that $c(X)\in\mathcal{C}$ if and only if $c(X)=u(X)g(X)$, for some $u(X)\in\mathbb{F}_{q}[X]$, $\deg(u(X))\leqslant n-r-1$. The roots of $g(X)$ (in its splitting field) are called the roots of the code $\mathcal{C}$. The commonly studied case is that of $g(x)$ having only simple roots (i.e., no repeated root). This is guaranteed, for example, when $\gcd(n,q)=1$.
@@ II-A Linear-feedback shift registers (LFSRs)
The following theorem summarizes results from [GorKlap12] and [Gol67]:
@@
    where $\alpha_{i}\in\mathbb{F}_{2^{d_{i}}}$ is a root of $M_{i}$ and $\gamma_{i}\in\mathbb{F}_{2^{d_{i}}}$ $1\leqslant i\leqslant e$, and $\operatorname{Tr}$ is the trace function from the splitting field of $g$ to $\mathbb{F}_{2}$. If $g$ is the minimal polynomial of the sequence, then the minimal period is equal to $\operatorname{lcm}\{\operatorname{ord}(g_{i}):1\leqslant i\leqslant e\}$.
@@ II-B Galois-mode LFSRs
We now introduce Galois-mode linear-feedback shift registers. We refer the reader to [GorKlap12], although our treatment differs slightly.
@@ II-C Character sums
Theorem 8 (Weil-Carlitz-Uchiyama bound, [CarlitzUchiyama57]).
@@ III Burst-Covering Codes
In case (2) is attained with equality, the resulting code is in fact perfect, i.e., simultaneously $b$-burst correcting and $b$-burst covering. To the best of our knowledge, only a single construction is known for such a code, with $q=2$ and $b=2$ (see [Etz01b]).
@@ IV Burst-Covering Radius of Binary Cyclic Codes
Without loss of generality, suppose $d_{e}=\min_{i}d_{i}$. Thanks to Equation (5), we know that any linear combination $(i,f)$ that yields the syndrome $(0,\dots,0,1)$, needs to satisfy $f(\alpha_{1})=\dots=f(\alpha_{e-1})=0$. Since $g_{1},\dots,g_{e-1}$ are pairwise coprime, $\prod_{i=1}^{e-1}g_{i}|f$, and so $\deg(f)\geqslant\sum_{i=1}^{e-1}d_{i}=\deg(g)-d_{e}$. We also note that $\operatorname{LC}(i,Xf)=\operatorname{LC}(i+1,f)$, so we may assume $i$ is set such that $X\nmid f$. But that the coefficients of the linear combination described by $f$ satisfy $c_{0}\neq 0$ and $c_{\deg(f)}\neq 0$, which means that at least $\deg(g)-d_{e}+1$ consecutive columns are needed.
Again by Equation (5), any linear combination of columns $(i,f)$ which is equal to 0 has to satisfy $\prod_{i=1}^{e}g_{i}|f$, thus if $f\neq 0$, $\deg(f)\geqslant\deg(g)$. So, any non-trivial linear combination of columns which is 0 requires at least $\deg(g)+1$ consecutive columns. This means that the first $\deg(g)$ columns of $H$ are linearly independent, and a basis for $\mathbb{F}_{2}^{r}$, which implies that a window of size $\deg(g)$ suffices. ∎
@@
A $b$-burst-correcting code with parameters $n,n-r]_{2}$ which attains this bound is called optimal (not to be confused with perfect as defined for bound (2)). Optimal codes have been constructed for different values of the parameters, and their existence in general was proven in [AbdMcEOdlTil86]. For example, in [ElsSho62] it is shown that taking a binary cyclic code with generator polynomial
@@ IV-A Pattern frequencies in LFSR sequences
From Equation (9), we can lower bound $N$ by
@@ IV-B Improved bounds on the burst-covering radius
Consider an $n,n-r]_{2}$ binary cyclic code, $\mathcal{C}$, with generator polynomial $g=\prod_{i=1}^{e}g_{i}$, where $g_{i}\in\mathbb{F}_{2}[X]$, for $1\leqslant i\leqslant e$, are distinct irreducible factors of degrees $d_{1}\leqslant d_{2}\leqslant\dots\leqslant d_{e}$, respectively, and $r=\sum_{i=1}^{e}d_{i}$. Denote the burst-covering radius of $\mathcal{C}$ by $b$. Then:
@@
    $b\leqslant r-\min_{J\subseteq\{1,\dots,e\},J\neq\emptyset}\left(\log_{2}(\operatorname{lcm}(\operatorname{ord}(g_{j}):j\in J))-\sum_{j\in J}d_{j}/2\right)$
@@
If $s\leqslant\min_{j\in J}d_{j}$, then by Theorem 25 the number $N$ of runs of $s$ consecutive zeros in such a sequence satisfies
@@
Thus, it suffices to ask that $s\leqslant\log_{2}(\operatorname{lcm}\{\operatorname{ord}(g_{j}):j\in J\})-\sum_{j\in J}d_{j}/2$ to guarantee that the sequence contains a run of $s$ zeros. Notice that for $J=\{1\}$, an LFSR sequence with minimal polynomial $g_{1}$ will never have more than $d_{1}-1$ zeros, so the minimum taken over $J$ is already less than $d_{1}=\min_{1\leqslant i\leqslant e}d_{i}$ and we do not need to include this condition.
3) Thanks to Equation (7), it suffices to show that any LFSR sequence with connection polynomial $g$ contains a run of at least $d_{1}-1$ zeros. Due to the primitivity assumption, an LFSR sequence with minimal connection polynomial $g_{i}$, for $i=1$ or $i=2$, is a PN sequence. Thus, it contains a run of $d_{i}-1\geqslant d_{1}-1$ zeros.
@@
Theorem 33 1) shows that the immediate lower bound on the covering radius from Theorem 19 cannot be attained unless the lowest-degree irreducible factor in the generator polynomial for the code is primitive. On the other hand, 3) shows that there are codes which attain this lower bound. It is interesting to note that the conditions in 3) are sufficient, but need not be necessary. In fact, it might be possible that some stronger version of Equation (8) would allow to remove the conditions on $\gcd(d_{1},d_{2})$ altogether, as we do not have any example of primitive polynomials $g_{1}$ and $g_{2}$ for which $b>r-d_{1}+1$.
@@ V Burst-Covering Radius of Long BCH Codes and Melas Codes
The burst-covering radius $b$ of $\operatorname{Melas}(m)$ satisfies
@@
The proof follows from Corollary 32 by noting that the generator polynomial for the code is $g=M_{1}\cdot\overleftarrow{M_{1}}$, where $\overleftarrow{h}=X^{\deg(h)}h(X^{-1})$. Thus, $e=d=1$ and $u_{1}=t_{1}=1$. ∎
@@
$m$ BCH Melas Upper Bound
6 9 10 10
7 11 11 11
8 12 12 13
9 13 14 14
10 14 15 16
11 16 16 17
We can compute the actual value of the burst-covering radius for small codes by using a computer program. For example, for $\operatorname{BCH}(2,m)$ and $\operatorname{Melas}(m)$, the results are given in Table I. We conjecture that the bound in Theorem 35 is essentially tight. Proving a lower bound requires a result like the following:
@@
The burst-covering radius $b$ of $\operatorname{BCH}(e,m)$ satisfies

### v2

@@ Abstract
We define and study burst-covering codes. We provide some general bounds connecting the code parameters with its burst-covering radius. We then provide stronger bounds on the burst-covering radius of cyclic codes, by employing linear-feedback shift-register (LFSR) sequences. For the case of BCH codes we prove a new bound on pattern frequencies in LFSR sequences, which is of independent interest. Using this tool, we can bound the covering-radius of binary primitive BCH codes and Melas codes. We then present an efficient burst-covering algorithm for cyclic codes. Finally, we present a bound on the critical exponent of linear cyclic codes based on the burst-covering radius.
@@ I Introduction
Error-correcting codes and their geometric counterparts, covering codes, have a long and rich history of research (e.g., [22, 10]). From a geometric perspective, while the former pack the space with error balls, the latter cover it. In their intersection lie perfect codes, that manage to tile the space with error balls.
Many families of codes have been studied both for the their error-correction capabilities, as well as their covering parameters. A partial list of those contains MDS codes, cyclic codes (including BCH, dual BCH, Melas and Zetterberg codes), Reed-Muller codes, and of course, perfect codes (for details, see for example [22, 10]).
However, one family of codes is conspicuously missing from this list – codes for burst errors. A $b$-burst error is an error pattern all of whose erroneous symbols are confined to a contiguous block of $b$ positions. From an error-correction perspective, such codes are motivated by the existence of bursty channels, that tend to group together erroneous positions. Burst-correcting codes have a history of research almost as long as that of error-correcting codes, starting with [3, 16, 14]. The burst-correcting capability of cyclic codes was studied in [23], where it was characterized in terms of the shortest feedback shift register generating certain sequences. Some cyclic burst-correcting codes are optimal, in the sense that they have the closest possible integer parameters to those dictated by the ball-packing bound, making them almost perfect. Among these we mention the cyclic codes of [1, 2], and the cyclic two-dimensional codes of [27, Construction A]. To the best of our knowledge, only one construction of perfect burst-correcting codes is known [15], for binary codes with burst length b=2b=2111There is a subtlety in the definitions of optimal and perfect burst-correcting codes: while optimal codes consider bursts cyclically, perfect codes do not.. These codes are therefore the only known burst-covering codes.
@@
Burst-covering codes do not only fill a void in the theory of binary codes, but may also find use in data storage scenarios. Many applications require the implementation of database queries whose return value is a linear combination of database items with coefficients supplied by the user. Private information retrieval (PIR) protocols [8], partial-sum queries [6] and more recently, certain machine-learning inference implementations [26] that are based on ideas from generalized covering radii [12, 13], among others, all employ such queries. Traditionally, covering codes are used to answer such linear queries. This relies on the fact that any column vector can be obtained as a linear combination of at most $R$ columns of a parity-check matrix for the code, where $R$ is the code’s covering radius. This guarantees a bounded access complexity, speeding up the computation of the answer to the linear query.
@@
However, in some media types, a significant component in the time to answer a query is not only the number of items that need to be accessed, but also their spatial location (e.g., the seek time in HDDs). Such systems perform best when accessing contiguous blocks of data, as opposed to scattered random-access patterns. This resulted in a surge of interest in codes that take into account access patterns, and prioritize access in contiguous blocks [29, 7, 30]. In the context of our work, as we will prove, the parity-check matrix $H$ of a linear $b$-burst-covering code satisfies the property that any column vector can be obtained as a linear combination of a window of at most $b$ consecutive columns of $H$. Thus, employing a burst-covering code can address locality issues in computing linear database queries.
@@
Our main contributions are as follows: We first define burst-covering codes and derive basic bounds on their parameters. We then focus on the study of the burst-covering radius of binary cyclic codes. This is motivated by the fact that cyclic codes have a rich structure, they contain some very useful code families (e.g., BCH codes), and the fact that we have almost perfect cyclic burst-correcting codes [1, 2]. As we later show, the study of cyclic burst-covering codes is closely related to the analysis of pattern frequencies in linear-feedback shift-register (LFSR) sequences. We employ classic known bounds for the number of occurrences of subwords in these sequences, but for the relevant case of BCH codes we show that these fail to provide significant results. We thus prove a new result on pattern frequencies which is meaningful for BCH codes, and is of independent interest for the study of LFSR sequences. We use this to bound the burst-covering radius of binary BCH codes, and the closely related Melas codes. We also show that our analysis naturally gives rise to an efficient covering algorithm for binary cyclic codes. Finally, we present a new bound on the critical exponent of binary cyclic codes based on the burst-covering radius of the dual code, and show that it can be significantly better than Kung’s bound.
@@
The paper is organized as follows. Section II gives the necessary notation and known results to be used later. In Section III we define general linear burst-covering codes, and provide some bounds on their parameters. We then to study cyclic burst-covering codes in Section IV, and further focus on the burst-covering radius of BCH codes in Section V. An efficient covering algorithm is described in Section VI. The connection to the critical exponent of linear codes is presented in Section VII. We conclude in Section VIII with a summary of the results and some open questions.
@@ II Preliminaries
Given a vector $v=(v_{0},\dots,v_{n-1})\in\mathbb{F}_{q}^{n}$, a cyclic shift of $v$ is the vector $(v_{n-1},v_{0},v_{1},\dots,v_{n-2})$. A linear code $\mathcal{C}$ is said to be cyclic if $c\in\mathcal{C}$ implies the cyclic shift of $c$ is also in $\mathcal{C}$. Denote $\mathbb{F}_{q}[X]$ the set of polynomials in the unknown $X$, with coefficients from $\mathbb{F}_{q}$. With any vector $v$ we associate the polynomial $v(X)=\sum_{i=0}^{n-1}v_{i}X^{i}$. It is well known [22], that an $n,n-r]_{q}$ cyclic code $\mathcal{C}$ is an ideal in the ring of polynomials $\mathbb{F}_{q}[X]/(X^{n}-1)$. There exists a unique generator polynomial, $g(X)\in\mathbb{F}_{q}[X]$, $\deg(g(X))=r$, such that $c(X)\in\mathcal{C}$ if and only if $c(X)=u(X)g(X)$, for some $u(X)\in\mathbb{F}_{q}[X]$, $\deg(u(X))\leqslant n-r-1$. The roots of $g(X)$ (in its splitting field) are called the roots of the code $\mathcal{C}$. The commonly studied case is that of $g(X)$ having only simple roots (i.e., no repeated root). This is guaranteed, for example, when $\gcd(n,q)=1$.
@@ II-A Linear-feedback shift registers (LFSRs)
The following theorem summarizes results from [18] and [17]:
@@
    where $\alpha_{i}\in\mathbb{F}_{2^{d_{i}}}$ is a root of $g_{i}$ and $\gamma_{i}\in\mathbb{F}_{2^{d_{i}}}$ $1\leqslant i\leqslant e$, and $\operatorname{Tr}$ is the trace function from the splitting field of $g$ to $\mathbb{F}_{2}$. If $g$ is the minimal polynomial of the sequence, then the minimal period is equal to $\operatorname{lcm}\{\operatorname{ord}(g_{i}):1\leqslant i\leqslant e\}$.
@@ II-B Galois-mode LFSRs
We now introduce Galois-mode linear-feedback shift registers. We refer the reader to [18], although our treatment differs slightly.
@@ II-C Character sums
Theorem 8 (Weil-Carlitz-Uchiyama bound, [5]).
@@ III Burst-Covering Codes
In case (2) is attained with equality, the resulting code is in fact perfect, i.e., simultaneously $b$-burst correcting and $b$-burst covering. To the best of our knowledge, only a single construction is known for such a code, with $q=2$ and $b=2$ (see [15]).
@@ IV Burst-Covering Radius of Binary Cyclic Codes
Without loss of generality, suppose $d_{e}=\min_{i}d_{i}$. Thanks to (5), we know that any linear combination $(i,f)$ that yields the syndrome $(0,\dots,0,1)$, needs to satisfy $f(\alpha_{1})=\dots=f(\alpha_{e-1})=0$. Since $g_{1},\dots,g_{e-1}$ are pairwise coprime, $\prod_{i=1}^{e-1}g_{i}|f$, and so $\deg(f)\geqslant\sum_{i=1}^{e-1}d_{i}=\deg(g)-d_{e}$. We also note that $\operatorname{LC}(i,Xf)=\operatorname{LC}(i+1,f)$, so we may assume $i$ is set such that $X\nmid f$. But that the coefficients of the linear combination described by $f$ satisfy $c_{0}\neq 0$ and $c_{\deg(f)}\neq 0$, which means that at least $\deg(g)-d_{e}+1$ consecutive columns are needed.
Again by (5), any linear combination of columns $(i,f)$ which is equal to 0 has to satisfy $\prod_{i=1}^{e}g_{i}|f$, thus if $f\neq 0$, $\deg(f)\geqslant\deg(g)$. So, any non-trivial linear combination of columns which is 0 requires at least $\deg(g)+1$ consecutive columns. This means that the first $\deg(g)$ columns of $H$ are linearly independent, and a basis for $\mathbb{F}_{2}^{r}$, which implies that a window of size $\deg(g)$ suffices. ∎
@@
A $b$-burst-correcting code with parameters $n,n-r]_{2}$ which attains this bound is called optimal (not to be confused with perfect as defined for bound (2)). Optimal codes have been constructed for different values of the parameters, and their existence in general was proven in [1]. For example, in [14] it is shown that taking a binary cyclic code with generator polynomial
@@ IV-A Pattern frequencies in LFSR sequences
From (9), we can lower bound $N$ by
@@ IV-B Improved bounds on the burst-covering radius
Consider an $n,n-r]_{2}$ binary cyclic code, $\mathcal{C}$, with generator polynomial $g=\prod_{i=1}^{e}g_{i}$, where $g_{i}\in\mathbb{F}_{2}[X]$, for $1\leqslant i\leqslant e$, are distinct irreducible factors of degrees $d_{1}\leqslant d_{2}\leqslant\dots\leqslant d_{e}$, respectively, and $r=\sum_{i=1}^{e}d_{i}$. Denote the burst-covering radius of $\mathcal{C}$ by $b$. Then, the burst-covering radius $b$ of $\mathcal{C}$ satisfies:
@@
    $b\leqslant r-\min_{J\subseteq\{1,\dots,e\},J\neq\emptyset}\varphi(J),$
    where
    $\varphi(J)=\begin{cases}d_{j}-1&\text{ if }J=\{j\}\text{ and }g_{j}\text{ is primitive}\\ \log_{2}(\operatorname{lcm}(\operatorname{ord}(g_{j}):j\in J))-\sum_{j\in J}d_{j}/2&\text{ otherwise}\end{cases}$
@@
If $J=\{j\}$ for some $j$ and $g_{j}$ is primitive, then the sequence with connection polynomial $g_{j}$ is a PN sequence, and its maximum run of zeros has length $d_{j}-1$.
Now suppose $J$ is not a singleton corresponding to a primitive polynomial. If $s\leqslant\min_{j\in J}d_{j}$, then by Theorem 25 the number $N$ of runs of $s$ consecutive zeros in such a sequence satisfies
@@
Thus, it suffices to ask that $s\leqslant\log_{2}(\operatorname{lcm}\{\operatorname{ord}(g_{j}):j\in J\})-\sum_{j\in J}d_{j}/2$ to guarantee that the sequence contains a run of $s$ zeros.
Notice that $\varphi(\{1\})<d_{1}=\min_{1\leqslant i\leqslant e}d_{i}$ and we do not need to include the condition $s\leqslant\min_{j\in J}d_{j}$ explicitly in the formula.
3) Thanks to (7), it suffices to show that any LFSR sequence with connection polynomial $g$ contains a run of at least $d_{1}-1$ zeros. Due to the primitivity assumption, an LFSR sequence with minimal connection polynomial $g_{i}$, for $i=1$ or $i=2$, is a PN sequence. Thus, it contains a run of $d_{i}-1\geqslant d_{1}-1$ zeros.
@@
Theorem 33 1) shows that the immediate lower bound on the covering radius from Theorem 19 cannot be attained unless the lowest-degree irreducible factor in the generator polynomial for the code is primitive. On the other hand, 3) shows that there are codes which attain this lower bound. It is interesting to note that the conditions in 3) are sufficient, but need not be necessary. In fact, it might be possible that some stronger version of (8) would allow to remove the conditions on $\gcd(d_{1},d_{2})$ altogether, as we do not have any example of primitive polynomials $g_{1}$ and $g_{2}$ for which $b>r-d_{1}+1$.
@@ V Burst-Covering Radius of Long BCH Codes and Melas Codes
The burst-covering radius $b$ of $\operatorname{Melas}(m)$, $m\geqslant 3$, satisfies
@@
When $m\geqslant 3$, we have that $\alpha$ and $\alpha^{-1}$ are not conjugates, and therefore their respective minimal polynomials $M_{1}$ and $\overleftarrow{M_{1}}$ are distinct, where $\overleftarrow{h}=X^{\deg(h)}h(X^{-1})$. The proof follows from Corollary 32 by noting that the generator polynomial for the code is $g=M_{1}\cdot\overleftarrow{M_{1}}$. Thus, $e=d=1$ and $u_{1}=t_{1}=1$. ∎
@@
$m$ BCH Melas Upper Bound $M_{1}(X)$
6 9 10 10 $X^{6}+X^{4}+X^{3}+X+1$
7 11 11 11 $X^{7}+X+1$
8 12 12 13 $X^{8}+X^{4}+X^{3}+X^{2}+1$
9 13 14 14 $X^{9}+X^{4}+1$
10 14 15 16 $X^{10}+X^{6}+X^{5}+X^{3}+X^{2}+X+1$
11 16 16 17 $X^{11}+X^{2}+1$
12 17 17 19 $X^{12}+X^{7}+X^{6}+X^{5}+X^{3}+X+1$
13 18 18 20 $X^{13}+X^{4}+X^{3}+X+1$
14 19 20 22 $X^{14}+X^{7}+X^{5}+X^{3}+1$
We can compute the actual value of the burst-covering radius for small codes by using a computer program. For example, for $\operatorname{BCH}(2,m)$ and $\operatorname{Melas}(m)$, the results are given in Table I. The burst-covering radius can depend on the primitive root $\alpha$ chosen to define the code, and thus we include the chosen $M_{1}(X)$.
We conjecture that the bound in Theorem 35 is essentially tight. Proving a lower bound requires a result like the following:
@@
The burst-covering radius $b$ of $\operatorname{BCH}(e,m)$, satisfies